# Provably Faster Gradient Descent via Long Steps

### v1

@@ 3.3 Certificates of Straightforwardness Proving Faster Rates in Table 1
First, we prove the claimed rates for the $t=2$ and $t=3$ stepsize patterns of Table 1 by presenting exact members of $\mathcal{S}_{h,0,\Delta}$. Then, to handle larger values of $t$, we present a simple rounding approach able to produce members of $\mathcal{S}_{h,\epsilon,\Delta}$, often with $\epsilon$ around machine precision $\leq 10^{-9}$. This approach produced rational-valued certificates proving the rest of the claimed convergence rates in Table 1.
The exact rational arithmetic verifying the correctness of all certificates $(\lambda,\gamma)$ was done in Mathematica 13.0.1.0. Note that the entries in these certificates for $t\geq 7$ are entirely computer-generated and lack real human insight. As an example for reference, the certificate for $t=7$ is included in the appendix. Larger certificates are impractical to include here. For example, our $t=127$ rate is certificate $(\lambda,\gamma)$ has $32640$ nonzero entries. Certificates for every pattern in Table 1 and exact verifying computations are available at github.com/bgrimmer/LongStepCertificates.
@@
For any $\eta\in(0,3)$, consider the selection of $(\lambda,\gamma)$ given by:
@@
Since this convex condition is linear in $\eta$, it suffices to check it at $\eta=0$ and $\eta=3$. Moreover, for any $\eta\in(0,3)$, note this matrix has exactly two zero eigenvalues with associated eigenvectors spanning $(1/2,1/2,1,0)$ and $(1/2,1/2,0,1)$. The second positive semidefiniteness condition amounts to checking an update to this matrix of size $\Delta$ remains positive semidefinite, namely
$\begin{bmatrix}\sum_{i=0}^{t-1}(h_{i}+\epsilon)&m_{h}(\gamma)^{T}\\ m_{h}(\gamma)&M_{h}(\lambda+\Delta\gamma)\end{bmatrix}=\begin{bmatrix}\frac{9-2\eta}{2}&\frac{-(3-\eta)}{2}&\frac{-(6-\eta)}{4}&\frac{-(6-\eta)}{4}\\ \frac{-(3-\eta)}{2}&\frac{1}{2}&\frac{2-\eta}{4}&\frac{2-\eta}{4}\\ \frac{-(6-\eta)}{4}&\frac{2-\eta}{4}&\frac{1}{2}&\frac{1}{2}\\ \frac{-(6-\eta)}{4}&\frac{2-\eta}{4}&\frac{1}{2}&\frac{1}{2}\end{bmatrix}+\begin{bmatrix}0&0&0&0\\ 0&\frac{-3}{2}&\frac{6-\eta}{4}&\frac{6-\eta}{4}\\ 0&\frac{6-\eta}{4}&0&0\\ 0&\frac{6-\eta}{4}&0&0\end{bmatrix}\Delta\succeq 0\ .$
One can check this added matrix term is positive semidefinite on the subspace spanned by $(1/2,1/2,1,0)$ and $(1/2,1/2,0,1)$ (again by checking when $\eta=0$ and $\eta=3$ and then using convexity). As a result, positive semidefiniteness is maintained for $\Delta$ small enough. Exact arithmetic verifying all of these claims are given in the associated Mathematica notebook. Hence $(\lambda,\gamma)\in\mathcal{S}_{h,0,\Delta}$, proving the main claim by Theorem 3.1 and the claimed convergence rate by Theorem 2.1. ∎
@@
The stepsize pattern $h=(1.5,4.9,1.5)$ is straightforward. Hence gradient descent (1.3) alternating between these three stepsizes has every $T=3s$ satisfy
@@
$\lambda=\begin{bmatrix}0&0&0&0&0\\ 0&0&1.95&0.003&0.007\\ 0&0.95&0&0.5&0.5\\ 0&0.006&0&0&0.51\\ 0&0.004&0&0.013&0\end{bmatrix}\ ,\qquad\gamma=\begin{bmatrix}0&0.005&7.825&3.9497&4.0203\\ 0&0&-5.24&-10.555&0\\ 0&0&0&7.9&-5.315\\ 0&0&0&0&1.2947\\ 0&0&0&0&0\end{bmatrix}\ .\ \qed$
@@
For larger settings $t\geq 7$, determining an exact member of $\mathcal{S}_{h,0,\Delta}$ proved difficult. So we resort to a fully automated construction of a convergence rate certificate by first numerically computing an approximate member of $\mathcal{S}_{h,0,\Delta}$ (via an interior point method) and then rounding to a nearby rational-valued exact member of $\mathcal{S}_{h,\epsilon,\Delta}$ for some small $\epsilon$. In light of our Theorem 2.1, such rounding only weakens the resulting guarantee’s coefficient from $\mathrm{avg}(h)$ to $\mathrm{avg}(h)-\epsilon$.
Computer Generation of Convergence Rate Proof Certificates. Given a pattern $h$ and $\Delta$,
@@
(iv) If so, $(\hat{\lambda},\hat{\gamma})\in\mathcal{S}_{h,\epsilon,\Delta}$, certifying a $\frac{LD^{2}}{(\mathrm{avg}(h)-\epsilon)T}+O(1/T^{2})$ convergence rate, for
@@
The patterns of lengths $t\in\{7,15,31,63,127\}$ in Table 1 are all $\epsilon$-straightforward for $\epsilon\in\{10^{-9},10^{-9},10^{-11},10^{-3},10^{-4}\}$ and $\Delta\in\{10^{-5},10^{-6},10^{-8},10^{-7},10^{-8}\}$. Hence the convergence rates claimed in Table 1 hold for each corresponding “long step” gradient descent method.
@@
Certificates $(\hat{\lambda},\hat{\gamma})\in\mathcal{S}_{h,\epsilon,\Delta}$ (produced via the above procedure) are available along with exact arithmetic validation at github.com/bgrimmer/LongStepCertificates. ∎
@@ 4 Future Directions
We have shown that using nonconstant, long stepsize patterns improves the performance guarantees of gradient descent. This runs contrary to widely held intuitions regarding constant stepsize selections and the importance of monotone objective decreases. Instead, we show that long-run performance improves by periodically taking (very) long steps that may increase the objective value in the short term. We accomplish this by developing a technique based on computer-generatable proof certificates to analyze the collective effect of many varied stepsizes at once. We conclude by discussing a few possible future improvements on and shortcomings of this technique.
@@
One practical drawback of the method (1.3) is the requirement that one knows $L$. Since our analysis in Theorem 2.1 only relies on decreases in objective value after $t$ steps, backtracking linesearch schemes or other adaptive ideas may be applicable. Such an approach could check a sufficient decrease condition every $t$ iterations, halving the estimate of $L$ and repeating those steps if not.
Future Improvements in Analysis Techniques. Almost all iterative optimization methods are analyzed based on deriving a recurrence, ensuring progress at every iteration. Our techniques deviate from this norm, only requiring improvement by a fixed horizon of length $t$. This deviation is also true of Young’s gradient descent method [19] for the strongly convex quadratic setting, which only ensures an optimal rate upon completion of the pattern. Determining which other first-order methods could benefit from collectively analyzing iterations up to a fixed horizon is an interesting open direction. Bregman, inexact, and stochastic variants of gradient descent are natural candidates.
Future works may improve our analysis by considering more sophisticated Lyapunov functions. Our proofs are only concerned with the eventual decrease of the objective gap. The analysis of optimal accelerated and subgradient methods relies additionally on the decreasing distance to a minimizer or a decreasing combination thereof. Identifying more tailored Lyapunovs and stepsize patterns guaranteed to eventually decrease it may lead to stronger guarantees.
Future Improvements in Computational Aspects. We observed that numerically computed primal optimal solutions to (3.2) were rank-one for all considered straightforward patterns. This corresponds to the worst-case objective function being essentially one-dimensional. This property was not used herein but could likely be leveraged to enable customized solvers for evaluating $p_{L,D}(\delta)$ and checking membership of $\mathcal{S}_{h,0,\Delta}$. Such improvements in tractability for SDPs with rank-one solutions have been studied widely [24, 25, 26, 27, 28, 29, 30] and may enable the search for longer, provably faster straightforward stepsize patterns than shown here.
As another avenue of improvement, note that any certificates produced by using floating point arithmetic followed by a rounding step (as done here) will likely lose a small $\epsilon$ amount in the rate. The use of an algebraic solver, like SPECTRA [31], could enable the automated production of exact certificates of $\epsilon=0-$straightforwardness as well as being able to certify when $\mathcal{S}_{h,0,\Delta}$ is empty.
@@ Acknowledgements.
The author thanks Fanghua Chen for preliminary discussions based on numerical observations that motivated this work. The openly released Performance Estimation Problem Branch-and-Bound solver and software of [12] was especially helpful in these explorations and for building intuitions. Additional thanks to Axel Böhm for providing useful feedback on the initial presentation of this work.

### v5

@@ 3.3 Certificates of Straightforwardness Proving Guarantees in Table 1
First, we prove the claimed guarantees for the $t=2$ and $t=3$ stepsize patterns of Table 1 by presenting exact members of $\mathcal{S}_{h,0,\Delta}$. Then, to handle larger values of $t$, we present a simple rounding approach able to produce members of $\mathcal{S}_{h,\epsilon,\Delta}$, often with $\epsilon$ around the accuracy of our SDP solves $\approx 10^{-9}$. This approach produced rational-valued certificates proving the rest of the claimed convergence guarantees in Table 1.
The exact rational arithmetic verifying the correctness of all certificates $(\lambda,\gamma)$ was done in Mathematica 13.0.1.0. Note that the entries in these certificates for $t\geq 7$ are entirely computer-generated and lack real human insight. As an example for reference, the certificate for $t=7$ is included in the appendix. Larger certificates are impractical to include here. For example, our $t=127$ guarantee is certificate $(\lambda,\gamma)$ has $32640$ nonzero entries. Certificates for every pattern in Table 1 and exact verifying computations are available at github.com/bgrimmer/LongStepCertificates.
@@
For any $\eta\in(0,3)$, consider the selection of $(\lambda,\gamma)$ given by
@@
Since this convex condition is linear in $\eta$, it suffices to check it at $\eta=0$ and $\eta=3$. Moreover, for any $\eta\in(0,3)$, note this matrix has exactly two zero eigenvalues with associated eigenvectors spanning $(1/2,1/2,1,0)$ and $(1/2,1/2,0,1)$. The second positive semidefiniteness condition amounts to checking an update to this matrix of size $\Delta$ remains positive semidefinite, namely $\begin{bmatrix}\sum_{i=0}^{t-1}(h_{i}+\epsilon)&m_{h}(\gamma)^{T}\\ m_{h}(\gamma)&M_{h}(\lambda+\Delta\gamma)\end{bmatrix}$ which equals
$\begin{bmatrix}\frac{9-2\eta}{2}&\frac{-(3-\eta)}{2}&\frac{-(6-\eta)}{4}&\frac{-(6-\eta)}{4}\\ \frac{-(3-\eta)}{2}&\frac{1}{2}&\frac{2-\eta}{4}&\frac{2-\eta}{4}\\ \frac{-(6-\eta)}{4}&\frac{2-\eta}{4}&\frac{1}{2}&\frac{1}{2}\\ \frac{-(6-\eta)}{4}&\frac{2-\eta}{4}&\frac{1}{2}&\frac{1}{2}\end{bmatrix}+\begin{bmatrix}0&0&0&0\\ 0&\frac{-3}{2}&\frac{6-\eta}{4}&\frac{6-\eta}{4}\\ 0&\frac{6-\eta}{4}&0&0\\ 0&\frac{6-\eta}{4}&0&0\end{bmatrix}\Delta$
must be positive semidefinite. One can check this added matrix term is positive semidefinite on the subspace spanned by $(1/2,1/2,1,0)$ and $(1/2,1/2,0,1)$ (again by checking when $\eta=0$ and $\eta=3$ and then using convexity). As a result, positive semidefiniteness is maintained for $\Delta$ small enough. Exact arithmetic verifying all of these claims are given in the associated Mathematica notebook. Hence $(\lambda,\gamma)\in\mathcal{S}_{h,0,\Delta}$, proving the main claim by Theorem 3.1 and the claimed convergence guarantee by Theorem 2.1. ∎
@@
The stepsize pattern $h=(1.5,4.9,1.5)$ is straightforward. Hence gradient descent (1.3) cycling through these three stepsizes has every $T=3s$ satisfy
@@
$\lambda=\begin{bmatrix}0&0&0&0&0\\ 0&0&1.95&0.003&0.007\\ 0&0.95&0&0.5&0.5\\ 0&0.006&0&0&0.51\\ 0&0.004&0&0.013&0\end{bmatrix},\ \gamma=\begin{bmatrix}0&0.005&7.825&3.9497&4.0203\\ 0&0&-5.24&-10.555&0\\ 0&0&0&7.9&-5.315\\ 0&0&0&0&1.2947\\ 0&0&0&0&0\end{bmatrix}.$
@@
For larger settings $t\geq 7$, determining an exact member of $\mathcal{S}_{h,0,\Delta}$ proved difficult. So we resort to a fully automated construction of a convergence guarantee certificate by first numerically computing an approximate member of $\mathcal{S}_{h,0,\Delta}$ (via an interior point method) and then rounding to a nearby rational-valued exact member of $\mathcal{S}_{h,\epsilon,\Delta}$ for some small $\epsilon$. In light of our Theorem 2.1, such rounding only weakens the resulting guarantee’s coefficient from $\mathrm{avg}(h)$ to $\mathrm{avg}(h)-\epsilon$.
Computer Generation of Convergence Proof Certificates. Given $h$ and $\Delta$,
@@
(iv) If so, $(\hat{\lambda},\hat{\gamma})\in\mathcal{S}_{h,\epsilon,\Delta}$, certifying $\frac{LD^{2}}{(\mathrm{avg}(h)-\epsilon)T}+O(1/T^{2})$ convergence, for
@@
The stepsize patterns of lengths $t\in\{7,15,31,63,127\}$ in Table 1 are all $\epsilon$-straightforward for $\epsilon\in\{10^{-9},10^{-9},10^{-11},10^{-3},10^{-4}\}$ with $\Delta\in\{10^{-5},10^{-6},10^{-8},10^{-7},10^{-8}\}$. Hence the convergence guarantees claimed in Table 1 hold for each corresponding “long step” gradient descent method.
@@
Certificates $(\hat{\lambda},\hat{\gamma})\in\mathcal{S}_{h,\epsilon,\Delta}$ (produced via the above procedure) with exact arithmetic validation are available at github.com/bgrimmer/LongStepCertificates. ∎
@@ 4 Future Directions
Building on existing performance estimation ideas, we have demonstrated an analysis technique capable of proving convergence of gradient descent using nonconstant, long stepsize patterns, which gain in strength as longer patterns are considered. This runs contrary to widely held intuitions regarding constant stepsize selections and the importance of monotone objective decreases. Instead, we show that long-run performance can improve by periodically taking (very) long steps that may increase the objective value in the short term. We accomplish this via computer-generated proof certificates bounding the effect of many steps collectively by providing solutions for a sequence of related performance estimation problems. We conclude by discussing a few possible future improvements on and shortcomings of this technique.
Possible Extension to More Families of Gradient Methods. Often, analysis techniques for gradient descent and its accelerated variants extend rather directly to constrained minimization or minimizing composite objectives $f(x)+r(x)$ by utilizing projections and proximal operators. PEP techniques naturally extend to these settings. Drori [35, Theorem 2.7 and 2.9] shows the tight convergence rate for projected gradient descent is strictly worse than the unconstrained setting, having rate $LD^{2}/4T$ instead of (1.2). A similar (small) worsening of the optimal rate also holds for accelerated proximal gradient methods [36]. This difficulty appears to extend to our straightforward analysis technique. Figure 2 shows the length two and three stepsize patterns that numerically satisfy a generalized notion of straightforwardness for gradient descent, projected gradient descent, and proximal gradient descent. The set of “straightforward” patterns for constrained is strictly smaller and composite yet strictly smaller. Consequently, strictly less aggressive stepsizes than Table 1 would be required for future projected or proximal extensions.
The PEP framework has previously been successfully employed in handling settings of inexact gradients [37] and relatively smooth optimization (via Bregman divergences) [7]. Determining the degree to which utilizing periodic long steps can improve the convergence of inexact, Bregman, and stochastic variations of gradient descent also provides interesting future directions.
Figure 2: For unconstrained, constrained, and composite minimization, the stepsize patterns satisfying a natural generalization of straightforwardness are shown. Patterns $(h_{1},h_{2})$ of length two and symmetric patterns $(h_{1},h_{2},h_{1})$ of length three were sampled at every $0.1$ increment. Straightforwardness was approximated by solving a performance estimation problem determining if $f(x_{t})-p_{\star}$ (or $f(x_{t})+r(x_{t})-p_{\star}$ for composite problems) is always at most $f(x_{0})-p_{\star}-\frac{\sum h_{i}}{LD^{2}}(f(x_{0})-p_{\star})^{2}$ (or $f(x_{0})+r(x_{0})-p_{\star}-\frac{\sum h_{i}}{LD^{2}}(f(x_{0})+r(x_{0})-p_{\star})^{2}$) given the initial gap was at most $\Delta=10^{-4}$.
@@
One practical drawback of the method (1.3) is the requirement that one knows $L$. Since our analysis in Theorem 2.1 only relies on decreases in objective value after $t$ steps, backtracking linesearch schemes or other adaptive ideas may be applicable. For example, one could consider an Armijo-type linesearching procedure, backtracking with an estimate of $\tilde{L}$ seeking $f(x_{t})\leq f(x_{0})-\frac{\sum h_{i}}{\tilde{L}}\|\nabla f(x_{0})\|^{2}$. Since convexity ensures $\|\nabla f(x_{0})\|^{2}\geq\delta_{0}^{2}/D^{2}$, this condition guarantees the descent required by straightforwardness. Analysis and practical development of such ideas are beyond our scope.
Future Improvements in Analysis Techniques. Future works may improve our analysis by considering other Lyapunov functions. The distance to optimal and norm of the gradient were both used in [4, Chapter 8]. Our proofs are only concerned with the eventual decrease of the objective gap. The analysis of optimal accelerated and subgradient methods relies additionally on the decreasing distance to a minimizer or a decreasing combination thereof. Identifying better Lyapunovs and stepsize patterns guaranteed to eventually decrease them may lead to stronger guarantees.
Future Improvements in Computational Aspects. We observed that numerically computed primal optimal solutions to (3.2) were rank-one for all considered straightforward patterns. This corresponds to the worst-case objective function being essentially one-dimensional. This is in line with many prior PEP-based analyses finding one-dimensional Huber functions often attain the worst-case performance. This property was not used herein but could likely be leveraged to enable customized solvers for evaluating $p_{L,D}(\delta)$ and checking membership of $\mathcal{S}_{h,0,\Delta}$. Such improvements in tractability for SDPs with rank-one solutions have been studied widely [38, 39, 40, 41, 42, 43, 44] and may enable the search for longer, provably faster straightforward stepsize patterns than shown here.
As another avenue of improvement, note that any certificates produced by using floating point arithmetic followed by a rounding step (as done here) will likely lose a small $\epsilon$ amount in the guarantee. The use of an algebraic solver, like SPECTRA [45], could enable the automated production of exact certificates of straightforwardness as well as being able to certify when $\mathcal{S}_{h,0,\Delta}$ is empty.
@@ Acknowledgements.
The author thanks Jason Altschuler for conversations at the Simons Insitute in Spring 2017 about the preliminary ideas subsequently developed in his excellent Master’s thesis work, Fanghua Chen for preliminary discussions based on numerical observations that motivated this work, and Adrien Taylor and Axel Böhm for providing useful feedback on the initial presentation of this work. The openly released Performance Estimation Problem Branch-and-Bound software of [13] was especially helpful in early explorations and building intuitions.